OpenStudy (anonymous):

Integral of .75x(1-x^2)dx from -1

OpenStudy (anonymous):

Wolfram keeps saying 1, my integrating keeps saying 0.

OpenStudy (anonymous):

So the integral is \[\int\limits_{-1}^{1}x[.75(1-x ^{2})]dx\]

OpenStudy (anonymous):

Which becomes... \[.75((x^{2}/2)-(x^{4}/4)) from -1 \to 1\]

OpenStudy (anonymous):

\[.75([((1)^{2}/2)-((1)^{4}/4)]-[((-1)^{2}/2)-((-1)^{4}/4)])\]

OpenStudy (anonymous):

\[.75[(1/4)-(1/4)]\]

OpenStudy (anonymous):

.75[0] = 0 What's wrong with my integral logic?

OpenStudy (anonymous):

you just use u sub u = 1-x^2 du= -2x dx

OpenStudy (amistre64):

\[.75x(1-x^2) = .75x-.75 x^3\]

OpenStudy (anonymous):

\[\frac{.75}{2} \int (u) \, du\]

OpenStudy (anonymous):

adjust the limit of integration accordingly

OpenStudy (amistre64):

\[\int .75x -.75x^3\ dx=\frac{.75}{2}x^2-\frac{.75}{4}x^4\]

OpenStudy (amistre64):

\[\frac{.75}{2}-\frac{.75}{4}-\frac{.75}{2}+\frac{.75}{4}=0\]

OpenStudy (amistre64):

unless stated otherwise, just let the integration take care of itself and the signed areas will cancel if they cancel

OpenStudy (anonymous):

\[\frac{1}{2} (-0.75) \int_0^0 u \, du\]

myininaya (myininaya):

\[\int\limits_{-1}^{1}\frac{3}{4}x(1-x^2) dx=\frac{3}{4} \int\limits_{-1}^{1}(x-x^3) dx\] \[\frac{3}{4}(\frac{x^2}{2}-\frac{x^4}{4})|_{-1}^{1}=\frac{3}{4}[(\frac{1}{2}-\frac{1}{4})-(\frac{1}{2}-\frac{1}{4})]\] \[\frac{3}{4}[0]=0\] and i like imran's way better less work

OpenStudy (anonymous):

wolfram say 0 too

OpenStudy (anonymous):

imram way is snappy. adjust the limits, forget the anti-derivative